

********************************************************************
UARS WEEKLY STATUS REPORT - 05 JAN 2001
********************************************************************
***********
SOLSTICE
***********
In support of this week's yaw maneuver, SOLSTICE operated in day only mode
as we were turned off and on each orbit to conserve power. This lasted for
two days during which we performed our normal solar observations but
gathered no stellar observations. The remainder of the week SOLSTICE
collected solar data on all available orbital tracking periods and stellar
calibration observations whenever possible. We continue to gather enough
data to meet all the SOLSTICE science objectives during this period of
reduced tape recorder operations.
Preliminary analysis of data following the operations anomaly in December
(reported in weekly report on 12/15/00) shows some degradation in the
SOLSTICE detectors. The instrument's gradual degradation is routinely
calibrated using data from the stellar observations. This one day step in
degradation will be included as part of this analysis as we have done in
the past for similar "one day hits". Two of the three SOLSTICE detectors
showed little degradation due to the exposure and have recovered nicely.
The mid-uv channel (170-315nm) showed the largest degradation of
approximately 40%. This degradation essentially decreases the lifetime of
the detector but we do not expect it to prematurely stop working.
We are planning a few special observations of Mercury to be carried out
starting next week when availabilities are favorable (Jan 8 - Jan 12)
***********
SUSIM
***********
SUSIM performed daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
AXIS 1 was activated on December 25 and deactivated on December 28. The
PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, unexplained missing transfers have
occurred to PEM2K::. PEM Telemetry is runing in duplicate streams through
PEM:: and PEM2K::. We find that there are occasional files which transfer
from the CDHF to PEM::, but are missing from PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the Monthly Status Review meeting for
UARS, held on 3 January 2001.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 9 February 2001.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 December 2000 (Orbit 50746, GMT Day 358) through
29 December 2000 (Orbit 50850, GMT Day 364).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 359/2339 50775 DAY ONLY
AXIS 359/2339 50775 AXIS 1 ON
PEM 360/1140 50783 DAY ONLY FIX
HALOE 361/2342 50805 POWER OFF
PEM 364/2345 50850 AXIS 1 OFF
The beta angle ranged between the angles of 25.5 degrees to 05.3 degrees
for this report period. The beta angle is now decreasing towards a
minimum angle of 2 degrees on 31 December 2000 (DOY 366).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.6 and 0.6 degrees C respectively.
Additional battery performance data for each battery on 22 December 2000
and 28 December 2000 are provided in Appendix D.
The clock error ranged from Ð10.9 to 10.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
¥ FOT = 11
¥ PEM = 03
¥ AXIS = 01
¥ HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
¥ FOT = 02
Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 00/363 show the observatory data loss to
be 300 hours, 01 minutes, 48.968 seconds. This shows an overall increase
of 24.080 seconds since the last report period This is a 0.39 percent
data loss, which equals a 99.61 percent data capture for the mission.
The decrease of 24.080 seconds occurred over one (1) day:
· 24.080 seconds on Day 363 due to a Late AOS
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week UARS successfully flew through the solar eclipse on 25
December 2000. The EOD SOC dropped 0.5 percent for one orbit and
immediately recovered on the next orbit.
The FOT and FSW group uplinked a one word patch to eliminate fixed point
overflow messages being received in the status buffer. The uplink was
successful and no overflow messages have been received since the patch
was made.
The FOT also generated the yaw plan and corresponding command sheets for
the upcoming yaw on 31 December 2000. All preparations have been made
and the load was built on 29 December 2000.
The performance of the TAMs continues to be monitored and analyzed on a
daily basis. Further testing of the TAMs on both the simulator and
onboard will continue next week. The FOT is trying to get the TAMs
ready before the next high beta period which will begin around 16
January 2000.
MEETINGS
The FOT attended a Year End Review Meeting on 27 December 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 31, 2000 (DOY 366).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 December 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
50750 358/0659 00-748 NBTR NOOPS D-358
50757 358/1818 00-750 NBTR NOOPS D-359
50764 359/0619 00-750 NBTR NOOPS D-359
50769 359/1319 00-750 NBTR NOOPS D-359
50772 359/1820 00-751 NBTR NOOPS D-360
50775 359/2339 00-752 PEM ZEPS/HEPS TO DAY ONLY
50775 359/2339 00-753 AXIS AXIS 1 ON
50780 360/0701 00-751 NBTR NOOPS D-360
50783 360/1140 00-754 PEM TO DAY ONLY FIX
50787 360/1822 00-755 NBTR NOOPS D-361
50794 361/0623 00-755 NBTR NOOPS D-361
50805 361/2342 00-757 HALOE POWER OFF
50811 362/0818 00-756 NBTR PLAYBACK
50832 363/1906 00-758 OBC UFLTR TLM PATCH
50845 364/1547 00-759 OBC/ACS Mag. Model time update
50850 364/2345 00-760 PEM AXIS 1 OFF
_______________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 December 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 December 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period
____________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 December 2000
22 Dec 2000 - GMT Day 357 Beta = 27.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +39.2/-28.0 +5.79/+4.83 +30.4/-14.4 24.5 1.040 21.3
3 +22.4/-112.0 +5.15/+3.87 +29.6/-15.2 24.5 1.040 21.0
28 Dec 2000 - GMT Day 363 Beta = 10.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------ ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-22.4 +4.51/+3.87 +29.6/-13.2 26.2 1.020 19.9
3 +22.4/-22.4 +3.23/+2.60 +28.8/-12.8 26.1 1.020 19.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ----------
no changes V/T 5
_______________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 December 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-057 FOT 50778 360/0439 14170 Generic Late cquisition
00-057 FOT 50827 363/1003 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
00-056 FOT 50584 347/0502 23302 Premature LOS.
00-057 FOT see list Dec 00 14170 Generic late Acquisition
SEVNTFW00/52